# algorithmic

### v2

@@ 5.5.2 Using Data More Efficiently
Semi-supervised training, blending labeled and unlabeled data, can leverage the unlabeled data to glean information about the task and the environment. This is most commonly done by generating pseudo-labels for the unlabeled data, either by leveraging model predictions or by making assumptions. The first approach is utilized by Cao et al. (2021b) and Zhan et al. (2021), which use generative models and GAN-based methods to mimic human preferences. Similarly, Park et al. (2022) expand their data set with high-confidence unlabeled samples based on the preference predictor’s evaluations. The second strategy, making assumptions to augment data, is showcased by Zhou & Xu (2020). They generate preference data by assuming that (i) human-written examples are better than model-written examples, (ii) human-written and model-written examples are indistinguishable amongst themselves, and (iii) generations of later model iterations are better than those of earlier ones.
@@ 5.5.3 Gathering Better Data
In addition to leveraging unlabeled data and using labels more efficiently, sample efficiency can be further increased by collecting more informative samples in the first place. This can either be achieved by selecting more informative samples from the experience buffer or by generating more informative experiences. While selecting informative samples from the experience buffer is addressed under active learning (see Section 4.1.1), this section focuses on generating more informative experiences.
@@ 6 Policy Learning
Figure 5: Highlighting the components of the RLHF loop discussed in this section.
After learning a reward model, or, more commonly, interleaved with reward model learning, the next step is to train a policy that maximizes the expected accumulated reward. This section will algorithms for policy learning, which can be categorized into two main techniques: adaptation of conventional RL algorithms and direct policy optimization (DPO).
@@ 6.1 Adaptation of RL Algorithms
Using the learned reward model, any standard RL algorithm (e.g., DQN, A3C, PPO, SAC) could potentially be applied to train a policy. However, in the setting of RLHF, this direct application may suffer from two issues: The non-stationary nature of the learned reward function in RLHF and the inaccuracy of intermediate reward models. We will discuss these issues and possible adaptations of RL algorithms to address them in the following.
@@
RL algorithms are designed to learn a policy that maximizes the expected accumulated reward in an MDP framework, which assumes a stationary reward function. The RLHF setting violates that assumption by periodically updating the reward model, leading to a non-stationary reward function.
@@
Later works have shown that value-based methods (possibly in an actor-critic scheme) can also be effective in RLHF (Ibarz et al., 2018; Lee et al., 2021b; Park et al., 2022; Liu et al., 2022; Xue et al., 2023b). One trick to make value-based methods work is to use the reward model to relabel the experiences in the replay buffer whenever it is updated (Ibarz et al., 2018; Lee et al., 2021b). Similar to conventional RL, the use of such a replay buffer can greatly decrease the amount of environment interactions necessary for successful learning. As demonstrated by Gulcehre et al. (2023), the sample efficiency can be increased even further by using offline-RL techniques in a growing-batch RL setting, an offline-RL technique that iteratively increases the size of the dataset by policy rollouts while still being more sample-efficient than online RL.
In addition to the basic RL approaches, there are also some policy learning approaches tailored specifically for RLHF. Wu et al. (2023) propose a policy gradient algorithm, called Pairwise Proximal Policy Optimization (P3O), as an alternative to PPO, which avoids estimating the value function and at the same time is provably invariant with respect to equivalent rewards (unlike PPO). In a similar vein, Zhu et al. (2023b) replace the KL-regularization of PPO by means of a squared error term of the logarithmic probabilities resulting in a seemingly more stable RL learner.
@@
Since the learned reward model, which is only an approximation of the true reward function, is used to train a policy, overoptimization (Gao et al., 2023) or reward hacking (Skalse et al., 2022) can happen. Section 5.2.5 discusses the interdependence of the reward model and the policy in more detail as well as possible improvements from the reward model side, while here we focus on how to adapt policy training to cope with possibly inaccurate rewards in general.
One approach is to regularize the policy so as not to diverge too much from human-given demonstrations. This is particularly common for language-model fine-tuning (Ouyang et al., 2022; Abramson et al., 2022), but Abramson et al. (2022) explores this for control as well. They found that this was important for some cases, in particular for deciding when to output language, but not for all. Going beyond KL-regularization, Moskovitz et al. (2024) investigate several techniques of constrained RL to only maximize rewards up to a threshold while avoiding excessive deviation from a pre-trained policy.
@@ 6.2 Framing RLHF for Generative Models as a Bandit Problem
So far, we have assumed that we ultimately want to solve a reinforcement learning problem represented by an MDP. However, especially with regard to the application of RLHF for the area of LLMs, there is now another simplified way of looking at the problem. Namely, as an instantiation of a (contextual) preference-based bandits problem (Bengs et al., 2021), which can of course be modeled by the more general case of a Markov decision process (MDP). In both cases, the focus is on the concept of tokens or rather sequences of tokens. However, in the MDP point of view, the state space ${\mathcal{S}}$ consists of all previous tokens and the prompt (represented as a sequence of tokens), while the action space ${\mathcal{A}}$ consists of all potential next tokens. A terminal state is often indicated here by the special token <eos> and trajectories are filled with this token until the maximum length $H$ is reached. Moreover, the transition function ${P}$ is degenerated (or deterministic) with being one only for the state that is the concatenation of the current state and the taken action. A (latent) reward is only received at the end of the trajectory giving rise to a sparse feedback scenario.
@@
Note that the bandit formulation considers an entire episode (response in the LLM context) as an action with a single associated reward, resulting in sparse feedback. This is in contrast to the standard RLHF formulation as it is often used in control settings, where it is assumed that the reward of a trajectory is composed of the sum of the rewards of individual steps, which allows the optimizer to distribute rewards densely as best fits the data. On an intuitive level, this leads to state-action pairs that often occur in preferred trajectories to be highly rewarded, without necessarily putting all reward on the terminal actions. In practice, this can lead to nicely-shaped reward functions (Christiano et al., 2017), which cannot directly be achieved in the bandit setting. However, Chan et al. (2024) show how to take advantage of the predominantly used transformer architecture for the reward model in order to obtain a denser reward, even when assuming the bandit setting: More specifically, since the transformer architecture maintains attention weights in the last layer for each token, these can be used to attribute the overall reward signal to individual tokens.
The main appeal of the bandit formulation is that it does not require exploration of the environment’s dynamics, since they are deterministic. It therefore enables simpler policy learning approaches, such as DPO (Rafailov et al., 2023) or $\Psi$PO (Azar et al., 2023), discussed in the following section.
@@ 6.3 Direct Policy Optimization
The two-phase approach involving utility learning and policy optimization is not the only viable path to learning a policy from human feedback. While we have previously discussed the case in which we learn a reward function from observed preferences by assuming a human feedback model, an emerging branch of the literature is concerned with circumventing the reward-learning step and using preferences directly to address the actual RL problem. Concrete approaches are DPO (Rafailov et al., 2023), SLiC-HF (Zhao et al., 2023), OPPO (Kang et al., 2023), DPPO (An et al., 2023), PRO (Song et al., 2024), RSO (Liu et al., 2024b), or by formulating policy search as a zeroth-order optimization (Tang et al., 2024). Azar et al. (2023) introduce an objective called $\Psi$-preference optimization ($\Psi$PO) that unifies the objective functions in DPO and RLHF. More specifically, for a specific instantiation of $\Psi$, the objective in $\Psi$PO recovers DPO and SLiC-HF. In addition, DPO has been further generalized to include diverse divergence constraints (Wang et al., 2024a). Besides, Hejna et al. (2024) propose contrastive preference learning based on a regret preference model instead of the usual one in RLHF. It is also possible to learn a $Q$ function from human preferences directly, which implies a policy without the need for separate policy- and reward-model training (Myers et al., 2023).
@@ 7 Theory
The field of RLHF has recently made some progress in terms of theoretical results, which we will discuss in this section. First, we consider the contributions where the goal is to learn a provable (near) optimal policy both in an online and offline fashion or even in a way that falls in between. Then, we discuss and highlight recent contributions related to different theoretical aspects of RLHF, such as its relation to the standard reward-based RL. Tables 6 and 7 provide a concise overview of the results for the online or offline policy learning setting. Here, $\mathcal{N}_{\mathcal{F}}(\epsilon,d)$ denotes the $\epsilon$-covering number of a set $\mathcal{F}$ under some metric $d$ [^4] The $\epsilon$-covering number is the minimum integer $N$ such that there exists a subset $\mathcal{F}^{\prime}\subset\mathcal{F}$ with $|\mathcal{F}^{\prime}|=N$, and for any $f\in\mathcal{F}$, there exists some $f^{\prime}\in\mathcal{F}^{\prime}$ satisfying $d(f,f^{\prime})\leq\epsilon$.. It is worth mentioning that (almost) all works have two standard assumptions, namely that the reward function is bounded and that the ground-truth reward, human feedback model, or transition dynamic is an element of the considered model space, respectively.
@@ 7.1 Policy Learning
In the literature focusing on theoretical results, there is a distinction (similar to the distinction made in standard RL) between an offline and online setting. In the former, learning is based on a given fixed data set, usually previously collected through an interaction with the environment. In contrast, in the online environment, one interacts directly with the environment to learn from real-time feedback and continuously updates one’s strategies based on the feedback received, allowing the agent to learn and adapt as it engages with the environment. Accordingly, an important component of the online variant is the sampling procedure, i.e., how the labels are selected. This is usually accomplished using an acquisition function that is based on uncertainty (see Section 4.1.1).
@@ 2. Unknown
Leveraging MaxInP; general function approximation General human feedback model class $\mathcal{F}_{\mathbb{T}}$ and general transition dynamic class $\mathcal{F}_{\mathbb{P}}$ with finite $l_{2}$-norm $\rho$-Eluder dimension $d_{\mathbb{T}}^{(2)}(\rho)$ and $d_{\mathbb{P}}^{(2)}(\rho)$, respectively High probability regret minimization w.r.t. optimal policy based on trajectory comparison feedback $\widetilde{\mathcal{O}}\left(\sqrt{d_{\mathbb{P}}(\tfrac{1}{T})HT\log\left(\mathcal{N}_{\mathcal{F}_{\mathbb{P}}}\left(\tfrac{1}{T},d\right)\right)}\right)\newline +\widetilde{\mathcal{O}}\left(\sqrt{d_{\mathbb{T}}(\tfrac{1}{T})T\log\left(\mathcal{N}_{\mathcal{F}_{\mathbb{T}}}\left(\tfrac{1}{T},d\right)\right)}\right)$
@@
$+\frac{H^{3}|\mathcal{S}|^{2}|\mathcal{A}|\log\left(\frac{H|\mathcal{S}||\mathcal{A}|}{\epsilon\delta}\right)}{\epsilon}\Bigg{)}$
@@
The first work dealing with the question of theoretical guarantees for learning an optimal policy from trajectory comparison feedback (see Section 3.2) in an online manner is by Novoseller et al. (2020). It laid the foundation for a paradigm subsequently embraced by many subsequent research endeavors: Adapting learning algorithms from the dueling or preference-based bandit literature (Yue & Joachims, 2009; Sui et al., 2018; Bengs et al., 2021) to the underlying situation with additional states. The preference-based bandit problem can be viewed as a preference-based RL problem with one state, so state transition dynamics must be considered accordingly for a fruitful adaptation. It is worth mentioning that Jain et al. (2015) used a quite similar idea before for feedback in the form of corrections (see Section 3.2) by resorting to the coactive learning setting (Shivaswamy & Joachims, 2012). Assuming the existence of a ground-truth context-trajectory scoring function and that the user’s feedback is informative, the Preference Perceptron algorithm by Shivaswamy & Joachims (2012) is used and analyzed in terms of its cumulative regret.
@@
Novoseller et al. (2020) suggest the Dueling Posterior Sampling (DPS), which is an adaptation of the self-sparring algorithm (Sui et al., 2017). It takes a Bayesian perspective on the problem and defines a Dirichlet prior on the dynamics and a Gaussian prior on the rewards that are subsequently updated, while the trajectories to be compared by the human labeler are chosen based on their (posterior) probability of being optimal [^5] The latter probability is assessed by posterior sampling; a commonly used technique in the bandit literature used by so-called Thompson Sampling strategies, see Lattimore & Szepesvári (2020) for more details.. Assuming a linear link function (see Section 5.1.4) as well as a tabular MDP, it is shown that DPS is (i) consistent, i.e., converges in distribution to the optimal policy, and (ii) achieves an asymptotic expected regret bound (see Table 6).
@@
Xu et al. (2020) combine dynamic programming and policy search with a black-box preference-based bandit algorithm for each state to design routines that return an almost optimal (a.k.a. $\varepsilon$-optimal) policy with high probability [^6] This is a so-called PAC learning setting (Valiant, 1984) in which the goal of finding the/an optimal object is relaxed to finding a “good enough” object, usually specified by some distance measure on the object domain. . The first routine, called Preference-based Policy Search (PPS), requires access to a simulator, while the second routine, called Preference-based Exploration and Policy Search (PEPS), gets rid of this requirement by exploring the state space by means of an auxiliary synthetic reward function. By assuming that the probability of one policy dominating another policy is bounded uniformly over all states from below by a multiplicative of their value function, they show generic upper bounds for both routines on the number of pairwise trajectory comparisons (see Table 6). If these dominance probabilities have even more structural properties, such as fulfilling stochastic transitivity or stochastic triangle inequality (see Haddenhorst et al. (2020); Bengs et al. (2021)), then these upper bounds can be further refined.
@@
A follow-up work by Saha et al. (2023) assumes a feature embedding of trajectories that gives rise to a feature embedding of policies and adapts the MaxInP algorithm (Saha, 2021) for contextual dueling bandits by essentially viewing the policy embeddings as the contexts. More precisely, assuming a logistic link function (see Section 5.1.4), confidence sets for the expected scores of the policies are constructed based on the maximum likelihood estimate (MLE), and the two policies with the highest uncertainty in terms of maximal variance are used to sample a trajectory, respectively, to be compared. In this way, the logistic preference-based reinforcement learning (LPbRL) is derived and also extended to the case of unknown dynamics by taking the uncertainty regarding the dynamics into account when constructing the confidence sets. For both cases, i.e., known or unknown dynamics, upper bounds on the regret of LPbRL are shown (see Table 6).
@@
In contrast to previous work that all considers tabular MDPs, Chen et al. (2022) consider the case of a general unknown human feedback model and unknown dynamics each from function classes with a finite Eluder dimension [^7] Roughly speaking, the Eluder dimension of a function class refers to the number of worst-case errors one must make to identify an unknown function from that class. (Russo & Van Roy, 2013). They propose and analyze the Preference-based Optimistic Planning (PbOP) algorithm, which essentially follows a similar design as LPbRL but uses least-square estimates for the human feedback model and transitions dynamics along with confidence sets based on them. Moreover, they derive lower bounds for the regret of any learning algorithm by reducing the once-per-episode-feedback RL problem (Chatterji et al., 2021) to the PbRL problem. Finally, they extend their analysis to the case of $K$-wise comparisons, where one obtains all $\binom{K}{2}$ pairwise comparisons for $K$ many queried trajectories. In essence, the regret term coming from the human feedback model class improves by a factor of $\sqrt{K}$.
@@
Wu & Sun (2024) consider a similar learning scenario as Saha et al. (2023) but with the additional objective to keep the number of queries of trajectory comparisons low, which is a combination of two competing objectives also studied in the bandit literature (Degenne et al., 2019). For this purpose, they suggest the Preference-based and Randomized Least-Squares Value Iteration (PR-LSVI) algorithm, which combines least-squares value iteration with a perturbed state-action-based reward model with Gaussian noise for regret minimization; a similar idea to CoLSTIM suggested for contextual dueling bandits (Bengs et al., 2022). More specifically, in each time step, the policy maximizing the value function of the perturbed state-action-based reward model and the policy maximizing the later in the previous time steps are “played”. By sampling trajectories of these two policies and computing their expected absolute reward difference (based on the perturbed state-action-based reward model) as a measure of uncertainty, preference feedback for these two trajectories is queried if the uncertainty exceeds a certain threshold. Moreover, they also suggest a posterior sampling counterpart of this algorithm, the Preference-based Thompson Sampling (PbTS) algorithm, and analyze it in terms of Bayesian quantities.
@@
Recently, Chakraborty et al. (2024) proposed a bilevel optimization problem that generalizes the standard optimization problem for RLHF with trajectory feedback and the negative log-likelihood as a loss function (i.e., (2)). This problem, which they call PARL (Policy Alignment in Reinforcement Learning), is characterized by explicitly taking into account the dependence on the data-collecting process at one level for the optimal policy parameters at the other level. For this problem, A-PARL is proposed, which is shown to have an $O(1/T)$ convergence rate under specific assumptions, where $T$ is the number of iterations.
@@
Finally, for the LLM training scenario, there are two perspectives on the problem (MDP vs. contextual dueling bandits) as mentioned in Section 6. Accordingly, work from the field of contextual preference-based bandits (Dudík et al., 2015; Saha, 2021; Saha & Krishnamurthy, 2022; Bengs et al., 2022; Sekhari et al., 2023) can also be viewed as theoretical contributions for the RLHF setting for the LLM application. This view is advocated in particular by Xiong et al. (2024), who view the LLM fine-tuning task as a reverse-KL regularized contextual bandit problem. Typically the context is chosen externally, but Mehta et al. (2023) consider the learning variant in which the learning agent chooses the context as well. This variant is referred to as active contextual dueling bandits.

### v3

@@ 5.5.2 Using Data More Efficiently
Semi-supervised training, blending labeled and unlabeled data, can leverage the unlabeled data to glean information about the task and the environment. This is most commonly done by generating pseudo-labels for the unlabeled data, either by leveraging model predictions or by making assumptions. The first approach is used by Cao et al. (2021b), using supervised learning methods to predict human preferences, and Zhan et al. (2021), employing a GAN-based approach to learn and mimic human preference patterns. Similarly, Park et al. (2022) expand their data set with high-confidence unlabeled samples based on the preference predictor’s evaluations. The second strategy, making assumptions to augment data, is showcased by Zhou & Xu (2020). They generate preference data by assuming that (i) human-written examples are better than model-written examples, (ii) human-written and model-written examples are indistinguishable amongst themselves, and (iii) generations of later model iterations are better than those of earlier ones. This is closely related to RLAIF, where a pretrained language model is used to gather preference feedback, effectively leveraging the knowledge about human preferences learned by the pretrained language model through self-supervised learning.
@@ 5.5.3 Gathering Better Data
In addition to leveraging unlabeled data and using labels more efficiently, sample efficiency can be further increased by collecting more informative samples in the first place. This can either be achieved by selecting more informative samples from the experience buffer or by generating more informative experiences. While selecting informative samples from the experience buffer is addressed under active learning (see Section˜4.1.1), this section focuses on generating more informative experiences.
@@ 6 Policy Learning
Figure 5: RLHF diagram highlighting components discussed in this section.
After learning a reward model, or, more commonly, interleaved with reward model learning, the next step is to train a policy that maximizes the expected accumulated reward. This section will discuss algorithms for policy learning, which can be categorized into two main techniques: adaptation of conventional RL algorithms and DPO.
@@ 6.1 Adaptation of RL Algorithms
Using the learned reward model, any standard RL algorithm (e.g., DQN, A3C, PPO, SAC) could potentially be applied to train a policy. However, in the setting of RLHF, this direct application may suffer from two issues: The non-stationary nature of the learned reward function in RLHF and its inaccuracy, especially in the earlier stages of training. We will discuss these issues and possible adaptations of RL algorithms to address them in the following.
@@
RL algorithms are designed to learn a policy that maximizes the expected accumulated reward in an MDP framework, which assumes a stationary reward function. The RLHF setting violates this assumption by periodically updating the reward model, leading to a non-stationary reward function.
@@
Later works have shown that value-based methods (possibly in an actor-critic scheme) can also be effective in RLHF (Ibarz et al., 2018; Lee et al., 2021b; Park et al., 2022; Liu et al., 2022; Xue et al., 2023). One trick to make value-based methods work is to use the reward model to relabel the experiences in the replay buffer whenever it is updated (Ibarz et al., 2018; Lee et al., 2021b). Similar to conventional RL, the use of such a replay buffer can greatly decrease the amount of environment interactions necessary for successful learning. As demonstrated by Gulcehre et al. (2023), the sample efficiency can be increased even further by using offline-RL techniques in a growing-batch RL setting, an offline-RL technique that iteratively increases the size of the dataset by policy rollouts while still being more sample-efficient than online RL.
In addition to the basic RL approaches, there are also some policy learning approaches tailored specifically for RLHF. Wu et al. (2023) propose a policy gradient algorithm, called Pairwise Proximal Policy Optimization (P3O), as an alternative to PPO, which avoids estimating the value function and at the same time is provably invariant with respect to equivalent rewards (unlike PPO). In a similar vein, Zhu et al. (2023b) replace the KL-regularization of PPO with a squared error term of the logarithmic probabilities, resulting in a seemingly more stable RL learner.
@@
Since the learned reward model, which is only an approximation of the true reward function, is used to train a policy, overoptimization (Gao et al., 2023) or reward hacking (Skalse et al., 2022) can happen. Section˜5.2.5 discusses the interdependence of the reward model and the policy in more detail as well as possible improvements from the reward model side, while here we focus on how to adapt policy training to cope with possibly inaccurate rewards in general.
One approach is to regularize the policy so as not to diverge too much from human-given demonstrations using KL-divergence regularization (Jaques et al., 2017). This is particularly common for language-model fine-tuning (Ouyang et al., 2022; Abramson et al., 2022), but Abramson et al. (2022) explores this for control as well. They found that this was important for some cases, in particular for deciding when to output language, but not for all. Going beyond KL-regularization, Moskovitz et al. (2024) investigate several techniques of constrained RL to only maximize rewards up to a threshold while avoiding excessive deviation from a pre-trained policy.
@@ 6.2 Framing RLHF for Generative Models as a Bandit Problem
So far, we have assumed that we ultimately want to solve a reinforcement learning problem represented by an MDP. However, especially with regard to the application of RLHF in the area of LLMs, there is now another simplified way of looking at the problem. Namely, as an instantiation of a (contextual) preference-based bandits problem (Bengs et al., 2021), which can of course be modeled by the more general case of a Markov decision process (MDP). In both cases, the focus is on the concept of tokens or rather sequences of tokens. However, in the MDP point of view, the state space $\mathcal{S}$} consists of all previous tokens and the prompt (represented as a sequence of tokens), while the action space $\mathcal{A}$} consists of all potential next tokens. A terminal state is often indicated here by the special token <eos> and trajectories are filled with this token until the maximum length $H$ is reached. Moreover, the transition function P{P} is degenerate (or deterministic), with a value of one only for the state that is the concatenation of the current state and the taken action. A (latent) reward is only received at the end of the trajectory giving rise to a sparse feedback scenario.
@@
Note that the bandit formulation considers an entire episode (response in the LLM context) as an action with a single associated reward, resulting in sparse feedback from an MDP viewpoint. This is in contrast to the standard RLHF formulation as it is often used in control settings, where it is assumed that the reward of a trajectory is composed of the sum of the rewards of individual steps, which allows the optimizer to distribute rewards densely as best fits the data. On an intuitive level, this leads to state-action pairs that often occur in preferred trajectories to be highly rewarded, without necessarily putting all reward on the terminal actions. In practice, this can lead to nicely-shaped reward functions (Christiano et al., 2017), which cannot directly be achieved in the bandit setting. However, Chan et al. (2024) show how to take advantage of the predominantly used transformer architecture for the reward model in order to obtain a denser reward, even when assuming the bandit setting: More specifically, since the transformer architecture maintains attention weights in the last layer for each token, these can be used to attribute the overall reward signal to individual tokens.
[…]
@@ 6.3 Direct Policy Optimization
The two-phase approach involving utility learning and policy optimization is not the only viable path to learning a policy from human feedback. While we have previously discussed the case in which we learn a reward function from observed preferences by assuming a human feedback model, an emerging branch of the literature circumvents the reward-learning step and uses preferences directly to address the actual RL problem. One important representative algorithm in this direction is DPO (Rafailov et al., 2023). The key trick in this method is to reparameterize the reward model as a function of its optimal policy, which allows the likelihood of the observed feedback (still using the Bradley-Terry model) to be rewritten as a function of the policy parameter. Thus, a policy can be directly trained from pairwise comparisons by minimizing the following negative log likelihood computed on a dataset $\mathcal{D}$ of triplets $(x,y_{w},y_{l})$ with $x$ being a context (e.g., prompt to LLM) and $y_{w},y_{l}$ being two outputs (e.g., generated answers) such that $y_{w}\succ y_{l}$:
$\mathcal{L}_{DPO}(\pi_{\theta},\pi_{\mathrm{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\mathrm{ref}}(y_{l}\mid x)}\right)\right]$ (14)
where $\sigma$ is the logistic function and $\beta$ is a hyperparameter controlling the deviation with respect to a reference policy $\pi_{\mathrm{ref}}$ (e.g., pretrained LLM). This loss amounts to increasing the probability of generating better outputs while decreasing that of generating worse ones with respect to the reference policy.
Many improvements, variations, or alternatives to DPO have been proposed. Since those methods are not the main focus of our survey, we refer the interested readers to their respective papers for more details, e.g.,
[…]
@@ 7 Theory
The field of RLHF has recently made some progress in terms of theoretical results, which we will discuss in this section. First, we consider the contributions where the goal is to learn a provably (near) optimal policy both in an online and offline fashion or even in a way that falls in between. Then, we discuss and highlight recent contributions related to different theoretical aspects of RLHF, such as its relation to the standard reward-based RL. Tables˜4 and 5 provide a concise overview of the results for the online or offline policy learning setting. Here, $\mathcal{N}_{\mathcal{F}}(\epsilon,d)$ denotes the $\epsilon$-covering number of a set $\mathcal{F}$ under some metric dd [^5] The $\epsilon$-covering number is the minimum integer $N$ such that there exists a subset $\mathcal{F}^{\prime}\subset\mathcal{F}$ with $|\mathcal{F}^{\prime}|=N$, and for any $f\in\mathcal{F}$, there exists some $f^{\prime}\in\mathcal{F}^{\prime}$ satisfying $d(f,f^{\prime})\leq\epsilon$.. It is worth mentioning that (almost) all works have two standard assumptions, namely that the reward function is bounded and that the ground-truth reward, human feedback model, or transition dynamics are elements of the considered model space, respectively.
@@ 7.1 Policy Learning
In the literature focusing on theoretical results, there is a distinction (similar to the distinction made in standard RL) between an offline and online setting. In the former, learning is based on a given fixed data set, usually previously collected through an interaction with the environment. In contrast, in the online environment, one interacts directly with the environment to learn from real-time feedback and continuously updates one’s strategies based on the feedback received, allowing the agent to learn and adapt as it engages with the environment. Accordingly, an important component of the online variant is the sampling procedure, i.e., how the labels are selected. This is usually accomplished using an acquisition function that is based on uncertainty (see Section˜4.1.1).
@@ 2. Unknown
Leveraging MaxInP; general function approximation General model classes $\mathcal{F}_{\mathbb{T}}$ for feedback and $\mathcal{F}_{\mathbb{P}}$ for transition dynamics with finite $l_{2}$-norm $\rho$-Eluder dimension $d_{\mathbb{T}}^{(2)}(\rho)$ and $d_{\mathbb{P}}^{(2)}(\rho)$. High probability regret minimization w.r.t. optimal policy based on trajectory comparison feedback $\widetilde{\mathcal{O}}\left(\sqrt{d_{\mathbb{P}}(\tfrac{1}{T})HT\log\left(\mathcal{N}_{\mathcal{F}_{\mathbb{P}}}\left(\tfrac{1}{T},d\right)\right)}\right)\newline +\widetilde{\mathcal{O}}\left(\sqrt{d_{\mathbb{T}}(\tfrac{1}{T})T\log\left(\mathcal{N}_{\mathcal{F}_{\mathbb{T}}}\left(\tfrac{1}{T},d\right)\right)}\right)$
@@
[…]
@@
The first work dealing with the question of theoretical guarantees for learning an optimal policy from trajectory comparison feedback (see Section˜3.2) in an online manner is by Novoseller et al. (2020). It laid the foundation for a paradigm embraced by many subsequent research endeavors: Adapting learning algorithms from the dueling or preference-based bandit literature (Yue & Joachims, 2009; Sui et al., 2018; Bengs et al., 2021) to the underlying situation with additional states. The preference-based bandit problem can be viewed as a preference-based RL problem with one state, so state transition dynamics must be considered accordingly for a fruitful adaptation. It is worth mentioning that Jain et al. (2015) used a quite similar idea before for feedback in the form of corrections (see Section˜3.2) by resorting to the coactive learning setting (Shivaswamy & Joachims, 2012). Assuming the existence of a ground-truth context-trajectory scoring function and that the user’s feedback is informative, the Preference Perceptron algorithm by Shivaswamy & Joachims (2012) is used and analyzed in terms of its cumulative regret.
@@
Novoseller et al. (2020) suggest the Dueling Posterior Sampling (DPS), which is an adaptation of the self-sparring algorithm (Sui et al., 2017). It takes a Bayesian perspective on the problem and defines a Dirichlet prior on the dynamics and a Gaussian prior on the rewards that are subsequently updated, while the trajectories to be compared by the human labeler are chosen based on their (posterior) probability of being optimal [^6] The latter probability is assessed by posterior sampling; a commonly used technique in the bandit literature used by so-called Thompson Sampling strategies, see Lattimore & Szepesvári (2020) for more details.. Assuming a linear link function (see Section˜5.1.4) as well as a tabular MDP, it is shown that DPS is (i) consistent, i.e., converges in distribution to the optimal policy, and (ii) achieves an asymptotic expected regret bound (see Table˜4).
@@
Xu et al. (2020) combine dynamic programming and policy search with a black-box preference-based bandit algorithm for each state to design routines that return an almost optimal (a.k.a. $\varepsilon$-optimal) policy with high probability [^7] This is a so-called PAC learning setting (Valiant, 1984) in which the goal of finding the/an optimal object is relaxed to finding a “good enough” object, usually specified by some distance measure on the object domain. . The first routine, called Preference-based Policy Search (PPS), requires access to a simulator, while the second routine, called Preference-based Exploration and Policy Search (PEPS), gets rid of this requirement by exploring the state space by means of an auxiliary synthetic reward function. By assuming that the probability of one policy dominating another policy is bounded uniformly over all states from below by a multiplicative of their value function, they show generic upper bounds for both routines on the number of pairwise trajectory comparisons (see Table˜4). If these dominance probabilities have even more structural properties, such as fulfilling stochastic transitivity or stochastic triangle inequality (see Haddenhorst et al. (2020); Bengs et al. (2021)), then these upper bounds can be further refined.
@@
A follow-up work by Saha et al. (2023) assumes a feature embedding of trajectories that gives rise to a feature embedding of policies and adapts the MaxInP algorithm (Saha, 2021) for contextual dueling bandits by essentially viewing the policy embeddings as the contexts. More precisely, assuming a logistic link function (see Section˜5.1.4), confidence sets for the expected scores of the policies are constructed based on the maximum likelihood estimate (MLE), and the two policies with the highest uncertainty in terms of maximal variance are used to sample a trajectory, respectively, to be compared. In this way, the logistic preference-based reinforcement learning (LPbRL) is derived and also extended to the case of unknown dynamics by taking the uncertainty regarding the dynamics into account when constructing the confidence sets. For both cases, i.e., known or unknown dynamics, upper bounds on the regret of LPbRL are shown (see Table˜4).
@@
In contrast to previous work that all considers tabular MDPs, Chen et al. (2022) consider the case of a general unknown human feedback model and unknown dynamics each from function classes with a finite Eluder dimension [^8] Roughly speaking, the Eluder dimension of a function class refers to the number of worst-case errors one must make to identify an unknown function from that class. (Russo & Van Roy, 2013). They propose and analyze the Preference-based Optimistic Planning (PbOP) algorithm, which essentially follows a similar design as LPbRL but uses least-square estimates for the human feedback model and transitions dynamics along with confidence sets based on them. Moreover, Chen et al. (2022) derive lower bounds for the regret of any learning algorithm by reducing the once-per-episode-feedback RL problem (Chatterji et al., 2021) to the PbRL problem. Finally, they extend their analysis to the case of $K$-wise comparisons, where one obtains all $\binom{K}{2}$ pairwise comparisons for $K$ many queried trajectories. In essence, the regret term coming from the human feedback model class improves by a factor of $\sqrt{K}$.
@@
Wu & Sun (2024) consider a similar learning scenario as Saha et al. (2023) but with the additional objective to keep the number of queries of trajectory comparisons low, which is a combination of two competing objectives also studied in the bandit literature (Degenne et al., 2019). For this purpose, they suggest the Preference-based and Randomized Least-Squares Value Iteration (PR-LSVI) algorithm, which combines least-squares value iteration with a perturbed state-action-based reward model with Gaussian noise for regret minimization; a similar idea to CoLSTIM suggested for contextual dueling bandits (Bengs et al., 2022). More specifically, in each time step, the policy maximizing the value function of the perturbed state-action-based reward model and the policy maximizing the latter in the previous time steps are “played”. By sampling trajectories of these two policies and computing their expected absolute reward difference (based on the perturbed state-action-based reward model) as a measure of uncertainty, preference feedback for these two trajectories is queried if the uncertainty exceeds a certain threshold. Moreover, they also suggest a posterior sampling counterpart of this algorithm, the Preference-based Thompson Sampling (PbTS) algorithm, and analyze it in terms of Bayesian quantities.
@@
Recently, Chakraborty et al. (2024a) proposed a bilevel optimization problem that generalizes the standard optimization problem for RLHF with trajectory feedback and the negative log-likelihood as a loss function (i.e., (2)). This problem, which they call PARL (Policy Alignment in Reinforcement Learning), is characterized by explicitly taking into account the dependence on the data-collecting process at one level for the optimal policy parameters at the other level. For this problem, A-PARL is proposed, which is shown to have an $O(1/T)$ convergence rate under specific assumptions, where $T$ is the number of iterations.
@@
Finally, for the LLM training scenario, there are two perspectives on the problem (MDP vs. contextual dueling bandits) as mentioned in Section˜6. Accordingly, work from the field of contextual preference-based bandits (Dudík et al., 2015; Saha, 2021; Saha & Krishnamurthy, 2022; Bengs et al., 2022; Sekhari et al., 2023) can also be viewed as theoretical contributions for the RLHF setting for the LLM application. This view is advocated in particular by Xiong et al. (2024), who view the LLM fine-tuning task as a reverse-KL regularized contextual bandit problem. Typically the context is chosen externally, but Mehta et al. (2023) consider the learning variant in which the learning agent chooses the context as well. This variant is referred to as active contextual dueling bandits.